OpenStudy (anonymous):

Solve 11 – x = 3. A. –8 B. 8 C. 14 D. –14 I got -8 because i subtracted 3 from each side is it right? and if it isn't can someone explain it please

OpenStudy (anonymous):

I mean -8

OpenStudy (unklerhaukus):

add x to both sides, and take away 3 from both sides

OpenStudy (anonymous):

11-x = 3 11=3+x 11-3 = x x = ?

OpenStudy (anonymous):

8..?

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

So it was just the negative that i got mixed up with i guess :) Thanks!